# “Medium” LMs of Code in the Era of LLMs: Lessons From StackOverflow

### v2

@@ 3. Approach
StackOverflow (SO) contains code in various languages with rich natural language information about the code. In this section we describe how we design SOBert, a BERT model trained on bimodal data from 15 million StackOverflow questions and answers.
3.1. Data and Preprocessing
@@
We downloaded the SO data dump published on 7 March 2022 (SO, 2021) which includes posts from 2008-2022. The Stack Overflow data dump is a collection of structured data that includes all information publicly available on the website. The data is made available for download periodically and includes information such as user profiles, questions, answers, comments, tags, and votes. The data is provided in XML format and is compressed into a set of files that can be downloaded and processed using a variety of tools and programming languages. Each file corresponds to a specific type of data, and contains all the relevant information for that type. We loaded these files into SQL database tables and specifically worked with the ‘Posts’ and ‘Comments’ tables. Using the entire corpus of Stack Overflow posts would have introduced a significant amount of low quality, unengaged content that may not be representative of posts software engineers typically rely on. We thus make a design choice to prioritize quality content that demonstrated some level of community engagement by filtering the answer posts to extract only those that meet the criteria of having a minimum of one up-vote and at least one accompanying comment. We then use this filter and extract both answer posts along with all their associated comments. This yielded 15 million answers and 39.5 million comments (median 2, mean 2.68 comments per post).
3.2. Data decisions
Devlin et al. (Devlin et al., 2018) and Tabassum et al. (Tabassum et al., 2020) tokenize the datasets for BERT and BERTOverflow respectively using a WordPiece vocabulary with 30,522 tokens. We preprocess the data for SOBert using a 50,000 token Byte-Pair Encoded (BPE) vocabulary, trained using SentencePiece (Kudo and Richardson, 2018). BPE sub-tokenizes words by greedily finding the largest tokens in its vocabulary that collectively reconstruct the word. While common words such as “the” tend to appear as a single token in the vocabulary, rare words may be broken up into smaller pieces. Since BPE uses a large vocabulary of single characters as its base vocabulary, it avoids treating almost any token as “unknown” or out-of-vocabulary, in contrast to WordPiece, which tends to improve results on downstream tasks (Karampatsis et al., 2020). We trained our tokenizer using a randomly selected subset of 10% of the entire dataset (ca. 2GB) due to memory limitations. The input text is first cleaned following standard data cleaning procedures such as first replacing URLs and email addresses with generic [URL] and [EMAIL] tokens to abstract away identifiable information.
We make sure to keep blocks of code, demarcated with <code> </code> tags, in the input data to preserve the complete structure of each post. This ensures that any code snippets or programming language syntax are not altered or misrepresented during processing or analysis. The preprocessed text is then input into the SentencePiece tokenizer for subword tokenization. A reserved token, <RS>, is placed between the answer and each of the comments to signal to the model where each comment begins. This tokenized representation is then passed through a stack of Transformer layers, described next.
3.3. Model Design
@@
Figure 4. Histogram showing length buckets of post+comments samples
@@
For the BERT base model the maximum length of the input sequence for the model is 512 tokens and each token is represented with a 768 dimensional vector. The choice of 2048 as the maximum length of the input sequence for SOBert was informed by an analysis of the text lengths of combined posts + comments that we use as inputs to our approach, shown in Figure 4. This shows that rather few posts are no more than 512 tokens long, which has conventionally been used as the maximum input length (Feng et al., 2020; Wang et al., 2021). A slight majority of the posts contain up to 1024 tokens but many more posts span between 1024-2048 tokens. As the cost of encoding long inputs grows quadratically in attention-based models, we choose to set the maximum length of inputs in our study to 2048 to balance training cost and fidelity to the task domain.
@@
Batch size refers to the number of training examples that are processed together in a single iteration during training. Larger batch sizes tend to lead to better language model performance (Kaplan et al., 2020) by providing more stable gradients for each model update, but they also require more memory and computing resources. We use gradient accumulation to simulate large batch sizes on our limited hardware, allowing us to process the training data in batches of 0.5 million tokens for both models. DeepMind proposed a series of “scaling laws” that can be used to determine how to best balance a given compute budget between training larger models and providing models more data (Hoffmann et al., 2022). Their conclusion was that models should be trained with ca. 20 times more tokens than they have parameters. While models trained this way won’t have converged (training longer would continue to yield better performance), the empirical observation is that this minimizes loss for a given compute budget. For instance, if one had the budget to train a 1 billion parameter model with 50B tokens, it would instead be better to train a significantly larger (e.g., 1.5B parameter) model with 30B tokens. The latter would achieve a lower loss in the same training time. This “rule” thus provides a useful lower bound for the training budget that should be allocated to a given model. We determined that, with our compute budget (see Section 3.4), the largest model that we would be able to train to this “optimum” within ca. two weeks time contained ca. 760M parameters. We therefore aimed to train this model with at least 15B tokens. Our actual training resources enabled us to provide nearly twice as many tokens as this minimum, which naturally improved its performance (Figure 5). [^5] We follow the conventional GPT-2 architecture sizes here. The typical “next-size” model would contain 1.4B parameters, which we estimate we could only train with 10 - 15B tokens, falling well short of the compute-optimum.
@@
We train our models using the Megatron-LM toolkit (Shoeybi et al., 2019). This toolkit was introduced for efficient parallel model training of LLMs with up to 8.3B parameters. It provides powerful support for multi-GPU training with optimized kernels for many components. It also incorporates many architectural best-practices. Shoeybi et al.(Shoeybi et al., 2019) showed that rearranging the order of the layer normalization and the residual connections is critical to enabling the scaling of the BERT-style models beyond 336M parameters. We use the same architecture configuration for SOBert.
Figure 5. Training and Validation Loss comparison of SOBBase and SOBLarge.
3.4. Training details
@@
The training was performed on four NVIDIA Quadro RTX 8000 GPUs, each with 48GB of memory, hosted in an on-premise server. To calculate the cost of training our models for those without access to this hardware, we base our estimates on quotes for NVIDIA’s A6000 GPU series, as these are currently widely available via cloud computing environments and have the same memory capacity as our GPUs. Nevertheless, A6000 GPUs outperform the Quadro RTX 8000 by ca. 80% (gpu, 2023b). We thus take the GPU hours required to train each of our models and multiply them by the standard rate of an A6000 in the cloud, which is approximately \$1 per hour (gpu, 2023a), and by the ratio 1/1.8 to compensate for the performance disparity. Training the SOBertLarge model required ca. 30 days on four GPUs for a total of ca. 2,880 GPU hours, which is equivalent to \$1600. The SOBertBase model required just 7 days ( ca. 672 GPU hours) to train, which amounts to \$374.
@@
4. Evaluation
@@
In this section we define the metrics we use to evaluate and compare the models and describe each of the four StackOverflow (SO) related downstream tasks, most of which come from prior research or Kaggle competitions. The four tasks we consider are question quality prediction (Annamoradnejad et al., 2022), closed question prediction(Kag, 2013), named entity recognition (NER) (Tabassum et al., 2020) and a novel task we formulate called obsoletion detection. We try to cover different types of downstream tasks with varying amounts of data available for fine-tuning. The above tasks, described in detail below, include a multi-class classification problem with about 60k samples, a multi-class classification problem with 140k samples, a token-level tagging task with 15k samples and a binary class classification with just 942 samples.
@@
Relevance of tasks to Software Engineering (SE) community: We have selected these four tasks based on their significance within the software engineering (SE) community and the availability of data for effective model performance comparisons. The first task involves predicting the quality of questions on platforms like StackOverflow. By identifying low-quality questions, this task contributes to a more efficient and informative knowledge-sharing environment. The second task predicts question closures, assisting moderators in managing platform content and ensuring its quality. Named Entity Recognition (NER), the third task, extracts crucial technical entities from text, benefiting various SE tasks such as code search by linking APIs and libraries to documentation. Finally, obsoletion detection, our fourth task, plays a vital role in highlighting outdated code snippets, promoting code health and security by preventing the use of vulnerable, deprecated code. These tasks collectively empower the development of improved tools for software developers, ultimately contributing to a more efficient and secure SE environment.
@@
Our evaluation of the models involves a comparative analysis with two general-purpose BERT models, namely BERTBase (Devlin et al., 2018) and BERTLarge (Devlin et al., 2018). In addition, we also assess the performance of the models against a domain-specific model called BERTOverflow (Tabassum et al., 2020) as well as with ChatGPT (cha, 2023), although we do not fine-tune the latter. In all cases, each model is trained with the same hyper-parameters, typically involving a batch size of 32 samples and a learning rate of 1e-5. Small variations from task to task, mainly in the input length, are noted below. We note that, although our model is capable of handling longer sequences, we set the maximum sequence length to be the same as that of the baseline models for all our experiments.
@@
SOBert was pre-trained using the Masked Language Modeling (MLM) objective. This means that during its pre-training phase, it was exposed to a wide range of text from Stack Overflow, which may or may not include some answers that appear in the downstream tasks. However, it did not have access to the labels for any of the downstream tasks. This reduces the risk of confounding effects related to data leakage in downstream tasks.
4.1. Metrics
Two of our tasks (NER and Obsoletion) have imbalanced datasets. An imbalanced dataset is one in which the distribution of classes is not uniform, meaning that one class may be over-represented or under-represented in comparison to the other classes. If the model always predicts “not-obsolete” or “O” the recall will be high, while a model that nearly never predicts “obsolete” or any of the other NER tags except when highly certain will achieve high precision. Metrics such as accuracy also tend to overvalue the performance on the majority class in imbalanced datasets. Class-weighted accuracy adjusts the contribution of each class to the overall accuracy score based on the inverse of the class frequency, giving more weight to the underrepresented classes. This offers a better indicator of a model’s true performance. The formula for weighted accuracy is:
$Acc_{W}=\sum_{i=1}^{N}\frac{n}{n_{i}}*Acc_{i}$
Where $n_{i}$ is the number of instances belonging to class $i$, $n$ is the total number of instances in the dataset, $N$ is the total number of classes, and $Acc_{i}$ is the accuracy of the model on class $i$.
Similarly, weighted recall is a performance metric used in classification tasks to measure the ability of a model to correctly identify all positive samples, taking into account the class imbalance. Recall is a measure of how well the model can identify positive samples from the total number of positive samples in the dataset. The formula for weighted recall incorporates class-weights as follows:
$Recall_{W}=\frac{\sum_{i=1}^{N}w_{i}*TruePositives_{i}}{\sum_{i=1}^{N}w_{i}*ActualPositives_{i}}$
Where $n$ is the number of classes, $w_{i}$ is the weight assigned to each class $i$, $\text{TruePositives}_{i}$ is the number of correctly predicted positive samples for class $i$, and $\text{ActualPositives}_{i}$ is the total number of actual positive samples for class $i$. The class weights are typically selected according to their inverse frequency, as in the weighted accuracy formula.
Neither $Precision$ nor $Recall$ fully captures a model’s performance. $F1-score$ takes both $Precision$ and $Recall$ into account and is their harmonic mean. Its class-weighted formula is:
$F1_{W}=\frac{\sum_{i=1}^{N}w_{i}*F1_{i}*ActualPositives_{i}}{\sum_{i=1}^{N}w_{i}*ActualPositives_{i}}$
In which $F1_{i}$ is the F1 score for class $i$.
4.2. Task 1: Named entity recognition (NER)
@@
Tabassum et al.(Tabassum et al., 2020) release a dataset for the extraction of software-related named entities from StackOverflow. It spans roughly 10 years (from September 2008 to March 2018) of question-answer threads in which each token in each sentence is manually annotated with close to 50 types of entities. The StackOverflow NER corpus has 9,352 train, 2,942 development and 3,115 test sentences. We fine-tuned all the models with a batch size of 16, maximum sequence length of 128 (based on the maximum length of the samples in the dataset) and learning rate of 1e-4.666Hyper-parameters are particularly different here because this task involves one label per token, so a batch of 16 samples has far more labels than in the other, sequence-level prediction tasks. We use a weighted cross entropy loss function as the dataset is imbalanced with many more ‘O’ (no label) tags than others. We were able to replicate the results of (Tabassum et al., 2020) within a 5%-10% margin for all reported metrics with this configuration on both their dev and test sets. The results of this task are discussed in the next section.
4.3. Task 2: Quality Tag prediction
We use another Kaggle competition dataset for predicting question quality on SO (Annamoradnejad et al., 2022) for this downstream task. This dataset features 60,000 SO questions from 2016-2020 classified into three categories based on their quality:
  • HQ: High-quality posts without a single edit.
  • LQ_EDIT: Low-quality posts with a negative score, and multiple community edits. However, they still remain open after those changes.
  • LQ_CLOSE: Low-quality posts that were closed by the community without a single edit.

### v3

@@ III Approach
SO contains code in various languages with rich natural language information about the code. In this section we describe how we design SOBert, a BERT model trained on bimodal data from 15 million SO questions and answers.
III-A Data and Preprocessing
@@
We downloaded the SO data dump published on 7 March 2022 [43] which includes posts from 2008-2022. The SO data dump is a collection of structured data that includes all information publicly available on the website. The data is made available for download periodically and includes information such as user profiles, questions, answers, comments, tags, and votes. The data is provided in XML format and is compressed into a set of files that can be downloaded and processed using a variety of tools and programming languages. Each file corresponds to a specific type of data, and contains all the relevant information for that type. We loaded these files into SQL database tables and specifically worked with the ‘Posts’ and ‘Comments’ tables. Using the entire corpus of SO posts would have introduced a significant amount of low quality, unengaged content that may not be representative of posts software engineers typically rely on. We thus make a design choice to prioritize quality content that demonstrated some level of community engagement by filtering the answer posts to extract only those that meet the criteria of having a minimum of one up-vote and at least one accompanying comment. We then use this filter and extract both answer posts along with all their associated comments. This yielded 15 million answers and 39.5 million comments (median 2, mean 2.68 comments per post).
@@
Figure 4: Histogram showing length buckets of post+comments samples
III-B Data decisions
Devlin et al. [4] and Tabassum et al. [15] tokenize the datasets for BERT and BERTOverflow respectively using a WordPiece vocabulary with 30,522 tokens. We preprocess the data for SOBert using a 50,000 token Byte-Pair Encoded (BPE) vocabulary, trained using SentencePiece [42]. BPE sub-tokenizes words by greedily finding the largest tokens in its vocabulary that collectively reconstruct the word. While common words such as “the” tend to appear as a single token in the vocabulary, rare words may be broken up into smaller pieces. Since BPE uses a large vocabulary of single characters as its base vocabulary, it avoids treating almost any token as “unknown” or out-of-vocabulary, in contrast to WordPiece, which tends to improve results on downstream tasks [44]. We trained our tokenizer using a randomly selected subset of 10% of the entire dataset (ca. 2GB) due to memory limitations. The input text is first cleaned following standard data cleaning procedures such as first replacing URLs and email addresses with generic [URL] and [EMAIL] tokens to abstract away identifiable information.
We use Beautiful Soup [45] to remove all HTML tags except for <code> tags. We make sure to keep blocks of code, demarcated with <code> </code> tags, in the input data to preserve the complete structure of each post. This ensures that any code snippets or programming language syntax are not altered or misrepresented during processing or analysis. The preprocessed text is then input into the SentencePiece tokenizer for subword tokenization. A reserved token, <RS>, is placed between the answer and each of the comments to signal to the model where each comment begins. This tokenized representation is then passed through a stack of Transformer layers, described next.
III-C Model Design
SOBert was pre-trained using the Masked Language Model (MLM) objective. This involves predicting masked tokens within a sentence, helping the model to learn the statistical properties of the language. The MLM objective can be represented by the following formula:
MLM $\text{MLM Loss}=-\sum_{i=1}^{N}\log P(w_{i}|\text{context})$
Where $N$ represents the total number of masked tokens, $w_{i}$ represents each masked token, and $P(w_{i}|\text{context})$ represents the probability of predicting the correct token given the surrounding context.
@@ III-C Model Design
For the BERT base model the maximum length of the input sequence for the model is 512 tokens and each token is represented with a 768 dimensional vector. The choice of 2048 as the maximum length of the input sequence for SOBert was informed by an analysis of the text lengths of combined posts + comments that we use as inputs to our approach, shown in Figure 4. This shows that rather few posts are no more than 512 tokens long, which has conventionally been used as the maximum input length [14, 13]. A slight majority of the posts contain up to 1024 tokens but many more posts span between 1024-2048 tokens. As the cost of encoding long inputs grows quadratically in attention-based models, we choose to set the maximum length of inputs in our study to 2048 to balance training cost and fidelity to the task domain.
@@
Batch size refers to the number of training examples that are processed together in a single iteration during training. Larger batch sizes tend to lead to better language model performance [16] by providing more stable gradients for each model update, but they also require more memory and computing resources. We use gradient accumulation to simulate large batch sizes on our limited hardware, allowing us to process the training data in batches of 0.5 million tokens for both models. DeepMind proposed a series of “scaling laws” that can be used to determine how to best balance a given compute budget between training larger models and providing models more data [6]. Their conclusion was that models should be trained with ca. 20 times more tokens than they have parameters. While models trained this way won’t have converged (training longer would continue to yield better performance), the empirical observation is that this minimizes loss for a given compute budget. For instance, if one had the budget to train a 1 billion parameter model with 50B tokens, it would instead be better to train a significantly larger (e.g., 1.5B parameter) model with 30B tokens. The latter would achieve a lower loss in the same training time. This “rule” thus provides a useful lower bound for the training budget that should be allocated to a given model. We determined that, with our compute budget (see Section III-D), the largest model that we would be able to train to this “optimum” within ca. two weeks time contained ca. 760M parameters. We therefore aimed to train this model with at least 15B tokens. Our actual training resources enabled us to provide nearly twice as many tokens as this minimum, which naturally improved its performance (Figure 5). [^2] We follow the conventional GPT-2 architecture sizes here. The typical “next-size” model would contain 1.4B parameters, which we estimate we could only train with 10 - 15B tokens, falling well short of the compute-optimum.
Figure 5: Training and Validation Loss comparison of SOBBase and SOBLarge.
@@
We train our models using the Megatron-LM toolkit [17]. This toolkit was introduced for efficient parallel model training of LLMs with up to 8.3B parameters. It provides powerful support for multi-GPU training with optimized kernels for many components. It also incorporates many architectural best-practices. Shoeybi et al.[17] showed that rearranging the order of the layer normalization and the residual connections is critical to enabling the scaling of the BERT-style models beyond 336M parameters. We use the same architecture configuration for SOBert.
III-D Training details
@@
The training was performed on four NVIDIA Quadro RTX 8000 GPUs, each with 48GB of memory, hosted in an on-premise server. To calculate the cost of training our models for those without access to this hardware, we base our estimates on quotes for NVIDIA’s A6000 GPU series, as these are currently widely available via cloud computing environments and have the same memory capacity as our GPUs. A6000 GPUs outperform the Quadro RTX 8000 by ca. 80% [46], so we take the GPU hours required to train each of our models and multiply them by the standard rate of an A6000 in the cloud, which is approximately \$1 per hour [47], and by the ratio 1/1.8 to compensate for the performance disparity. Training the SOBertLarge model required ca. 30 days on four GPUs for a total of ca. 2,880 GPU hours at an A6000-equivalent cost of \$1600. The SOBertBase model required just 7 days ( ca. 672 GPU hours) to train, which amounts to \$374.
@@
III-E Limitations
While our assessment of SOBert includes a comprehensive analysis of four downstream tasks from well-established sources, it is important to recognize that the applicability of our findings to different tasks may differ. Given that SOBert is a predictive model, our focus is primarily on classification tasks rather than generative ones. Although we would have preferred to test baselines against multiple BERT models pre-trained on SO data such as SeBERT and PTM4Tag, resource limitations compelled us to select only the state-of-the-art BERTOverflow and two general BERT models of varying sizes. We discuss our selection rationale in detail in Section II. Additionally, the nature of API access from OpenAI implies that the performance of the language model can fluctuate over time due to model updates without prior notification, potentially affecting its performance. However, the fact that the LLMs are much less accurate in each task implies that it is unlikely to be able to bridge the absence of task specific training with examples alone.
IV Evaluation
@@
In this section, we establish the metrics employed for evaluating and contrasting the models, while providing a comprehensive overview of the four SO related downstream tasks. SO is known to cover numerous SE topics and attracts numerous software developers. Among these, three tasks have been sourced from prior research or Kaggle competitions, namely question quality prediction [19], closed question prediction [20], and named entity recognition (NER) [15]. Additionally, we introduce a novel task we have formulated, known as obsoletion detection. We try to cover different types of downstream tasks with varying amounts of data available for fine-tuning. The above tasks, described in detail below, include a multi-class classification problem with about 60k samples, a multi-class classification problem with 140k samples, a token-level tagging task with 15k samples. The new task, obsoletion detection, is a binary class classification task with just 942 samples.
@@
Selection Rationale for downstream tasks: Despite some claims that SO is losing relevance, the demand for human-generated responses to novel programming inquiries will persist and cannot be entirely supplanted by current technological advancements [48]. We have selected these four tasks based on their significance within the SE community, their relevance to the SO data source we used for training, and the availability of data for effective model performance comparisons. The first task involves predicting the quality of questions on platforms like SO. By identifying low-quality questions, this task contributes to a more efficient and informative knowledge-sharing environment. The second task predicts question closures, assisting moderators in managing platform content and ensuring its quality. NER, the first task, extracts crucial technical entities from text, benefiting various SE tasks such as code search by linking APIs and libraries to documentation. Finally, obsoletion detection, our fourth task, plays a vital role in highlighting outdated code snippets, promoting code health and security by preventing the use of vulnerable, deprecated code. These tasks are also relevant for evaluating the quality and safety of code generated by LLMs, which are increasingly being adopted by the SE community. These tasks collectively empower the development of improved tools for software developers, ultimately contributing to a more efficient and secure SE environment.
@@
Fine-tuning details: Our evaluation of the models involves a comparative analysis with two general-purpose BERT models, namely BERTBase [4] and BERTLarge [4]. In addition, we also assess the performance of the models against a domain-specific model called BERTOverflow [15]. In all cases, each model is trained with the same hyper-parameters, typically involving a batch size of 32 samples and a learning rate of 1e-5. Small variations from task to task, mainly in the input length, are noted below. We note that, although our model is capable of handling longer sequences, we set the maximum sequence length to be the same as that of the baseline models for all our experiments.
We also compare SOBert with LLMs from two model families for the obsoletion detection task. These include GPT-4 and GPT-3.5 as closed-source LLMs. We choose CodeLlama and StackLlama as open-source LLMs. We use zero-shot, few shot with CoT prompting for the closed-source LLMs. We fine-tuned the open-source LLMs using Quantized LoRA (QLoRA) and the Parameter-Efficient Fine-Tuning (PEFT) library. QLoRA involves training only a small set of quantized adapter weights instead of the full model weights, significantly reducing memory and computational requirements. We utilized the LoraConfig from the PEFT library to specify the QLoRA configuration and applied techniques like gradient checkpointing and mixed-precision training for efficient fine-tuning. The fine-tuning process was carried out using the SFTTrainer from the Transformers library. After fine-tuning, we merged the trained QLoRA adapter weights with the original CodeLLaMA model weights using PEFT’s merge_and_unload method, creating a single model file for inference. This approach enabled efficient adaptation of the large CodeLLaMA model while preserving its pre-trained knowledge.
@@
Data leakage: During the pre-training phase, SOBert was exposed to a wide range of text from SO, which may or may not include some answers that appear in the downstream tasks. However, it did not have access to the labels for any of the downstream tasks. This reduces the risk of confounding effects related to data leakage in downstream tasks.
IV-A Task 1: [Uncaptioned image] Named entity recognition (NER) [15]
@@
Tabassum et al.[15] release a dataset for the extraction of software-related named entities from StackOverflow. It spans roughly 10 years (from September 2008 to March 2018) of question-answer threads in which each token in each sentence is manually annotated with close to 50 types of entities. The StackOverflow NER corpus has 9,352 train, 2,942 development and 3,115 test sentences. We fine-tuned all the models with a batch size of 16, maximum sequence length of 128 (based on the maximum length of the samples in the dataset) and learning rate of 1e-4.333Hyper-parameters are particularly different here because this task involves one label per token, so a batch of 16 samples has far more labels than in the other, sequence-level prediction tasks. We use a weighted cross entropy loss function as the dataset is imbalanced with many more ‘O’ (no label) tags than others.
We were able to replicate the results of [15] within a 5%-10% margin for all reported metrics with this configuration on both their dev and test sets. The results of this task are discussed in the next section.
IV-B Task 2: [Uncaptioned image] Quality Tag prediction [19]
We utilize a Kaggle competition dataset focused on predicting question quality on SO for content moderation in our downstream task [19]. This dataset features 60,000 SO questions from 2016-2020 classified into three categories based on their quality - HQ: High-quality posts without a single edit, LQ_EDIT: Low-quality posts with a negative score, and multiple community edits. However, they still remain open after those changes, LQ_CLOSE: Low-quality posts that were closed by the community without a single edit.